OpenStudy (anonymous):

Prove that the lim x ->0 xcos(1/x) = 0 Prove that the lim x ->0 xcos(1/x) = 0 @Mathematics

OpenStudy (anonymous):

By the squeeze theorem\[0\le \lim_{x \rightarrow 0}\cos(1/x)\le1\]\[\lim_{x \rightarrow 0} (x \times 0)\le \lim_{x \rightarrow 0}xcos(1/x)\le \lim_{x \rightarrow 0}x\]\[0\le \lim_{x \rightarrow 0}xcos(1/x) \le 0\]therefore\[\lim_{x \rightarrow 0} xcos(1/x) = 0\]

OpenStudy (anonymous):

thnks I used -1 <= cos(1/x) <=1 -x <= xcos(1/x) <=x lim (-x) = 0 and lim x = 0

OpenStudy (anonymous):

o whoops yea that should be a -1

OpenStudy (anonymous):

cool but the squueze thm is the way to go thnks for confirming,